OpenStudy (swag):

inverse of the function y = 3x + 7

OpenStudy (anonymous):

y = 3x + 7 y-7 = 3x (y-7)/3 = x // just move x to the left x = (y-7)/3 replace all y's for x's and all x's for y's

OpenStudy (anonymous):

*replace the x with y^-1 and replace the y with x  After, make the y^-1 subject